A coffee merchant has coffee beans that sell for $9 per
pound and $12 per pound. The two types are to be mixed to create 100 lb of a mixture
that will sell for $11.25 per pound. How much of each type of bean should be
used in the mixture?
61. Business and finance. Peanuts are selling for $2 per pound, and cashews are sell

Answer by stanbon(48558)

A coffee merchant has coffee beans that sell for $9 per pound and $12 per pound. The two types are to be mixed to create 100 lb of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?
-------------
Let amt. of $9 beans be "x lbs ; Value of these is 9x dollars
------------------
Amt of $12 beans is "100-x" lbs ; Value of these is 12(100-x)=1200-12x dollars
---------------------
Amt of mixture is 100 lbs.; value is 11.25*100 = 1125 dollars
-----------------------
EQUATION:
9x + 1200-12x = 1125
-3x = -75
x = 25 lbs (amount of $9 beans in the mixture)
100-x = 75 lbs ( amount of $12 beans in the mixture)
================
Cheers,
Stan H.







61. Business and finance. Peanuts are selling for $2 per pound, and cashews are sell